# Interpolation theorems for conjugations and applications

### v1

@@ 3.1. Main results
$Vp(N,N^{*})y_{i}=p(-N,-N^{*})^{*}Vy_{i}=p(-N,-N^{*})^{*}x_{i}.$
This implies that if $L_{0}={\rm span}\big{\{}p(N,N^{*})x_{i}+q(N,N^{*})y_{i}:p,q\in\mathbb{C}[x,y],i\in I\big{\}}$, then $VL_{0}=L_{0}$. Taking the closure of $L_{0}$ in $H$, we obtain $VL_{N,Z}=L_{N,Z}$; moreover, since $V$ is anti-unitary, we get $VL_{N,Z}^{\perp}=L_{N,Z}^{\perp}$.
Let $V_{1}$ and $N_{1}$ denote the restrictions of $V$ and $N$ to $L_{N,Z}^{\perp}$, respectively. Then, $V_{1}$ is anti-unitary on $L_{N,Z}^{\perp}$, and the relation $VN=-N^{*}V$ implies $V_{1}N_{1}=-N_{1}^{*}V_{1}$. Consequently, $N_{1}$ is skew-symmetric by Lemma 3.5.
Now, by Lemma 2.11, for every Borel subset $\Delta\subseteq\mathbb{C}$, we have $VE_{N}(-\Delta)=E_{N}(\Delta)V$. Hence, using (2.9), for all $i,j\in I$,
$\langle E_{N}(\Delta)x_{i},y_{j}\rangle$ $=\langle E_{N}(\Delta)Vy_{i},Vx_{j}\rangle=\langle VE_{N}(-\Delta)y_{i},Vx_{j}\rangle$
$=\langle V^{-1}Vx_{j},E_{N}(-\Delta)y_{i}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle,$
and similarly, $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{N}(-\Delta)y_{j},y_{i}\rangle$.
(iii)$\implies$(i). Let $M=-N$. For every Borel subset $\Delta\subseteq\mathbb{C}$ and for all $i,j\in I$,
$\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle=\langle E_{-N}(\Delta)x_{j},y_{i}\rangle=\langle E_{M}(\Delta)x_{j},y_{i}\rangle,$
and similarly, $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{M}(\Delta)y_{j},y_{i}\rangle$. For convenience, put $H_{0}=L_{N,Z}$, $H_{1}=H_{0}^{\perp}$, $N_{0}=N_{|H_{0}}$ and $N_{1}=N_{|H_{1}}$. By Theorem 2.10, there exists a conjugation $C_{0}$ on $H_{0}$ such that $C_{0}N_{0}C_{0}=-N_{0}^{*}$ and $C_{0}x_{i}=y_{i}$ for every $i\in I$. Since $N_{1}$ is skew-symmetric, there exists a conjugation $C_{1}$ on $H_{1}$ such that $C_{1}N_{1}C_{1}=-N_{1}^{*}$, and thus the conjugation $C=C_{0}\oplus C_{1}$ satisfies all required properties. ∎
@@
(ii)$\implies$(i). Let $N_{0}$ and $N_{1}$ be the normal operators in the proof of Theorem 3.2. In the light of the same theorem, we only need to show that $N_{1}$ is skew-symmetric. This follows immediately from Theorem 3.8 and the facts that $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$ and $N=N_{0}\oplus N_{1}$ is skew-symmetric.
@@
$H_{0}=\bigvee\left(\sum_{i\in I}L_{N,\{x_{i}\}}+L_{N,\{y_{i}\}}\right)=\bigvee\left(\sum_{i\in I}L_{N_{0},\{x_{i}\}}+L_{N_{0},\{y_{i}\}}\right)$
@@
and the restrictions of $N_{0}$ to $L_{N_{0},\{x_{i}\}}$ and $L_{N_{0},\{y_{i}\}}$ are star-cyclic, Lemma 3.9 implies that $m({N_{0}}_{|{\rm ker}(N_{0})^{\perp}})<\infty$. This concludes the proof. ∎
4. Applications to some equations in $L^{\infty}(\mu)$
Throughout this section, let $\mu$ denote a finite measure on Borel subsets of $\mathbb{C}$ with compact support, and let $n$ be a positive integer. All functions are considered defined on $\mathbb{C}$. For a given function $f$, the notation $\widetilde{f}$ stands for the function defined by $\widetilde{f}(z)=\overline{f(-z)}$ for $z\in\mathbb{C}$.
Let $[\phi_{i,j}]_{1\leq i,j\leq n}$ be a matrix of functions in $L^{\infty}(\mu)$. We denote $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$ if the matrix $[\phi_{i,j}(z)]_{1\leq i,j\leq n}$ is unitary for $\mu$-almost every $z\in\mathbb{C}$. If, in addition, the functions $\phi_{i,j}$ satisfy the condition $\phi_{i,j}(z)=\phi_{j,i}(-z)$ for $\mu$-almost every $z\in\mathbb{C}$, then we write $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{SU}(\mu,n)$.
For a function $\phi\in L^{\infty}(\mu)$, $M_{\phi}$ denotes the multiplication operator by $\phi$ on $L^{2}(\mu)$; that is, $[M_{\phi}f](z)=\phi(z)f(z)$ for every $f\in L^{2}(\mu)$.
Let $\{f_{i}\}_{1\leq i\leq n}$ and $\{g_{i}\}_{1\leq i\leq n}$ be functions in $L^{2}(\mu)$ so that $\sum_{k=1}^{n}|f_{k}|^{2}=\sum_{k=1}^{n}|g_{k}|^{2}$ $\mu$-almost everywhere ($\mu$-a.e.). It is clear that we can find functions $\phi_{i,j}$ so that, for $\mu$-almost every $z\in\mathbb{C}$, the matrix $[\phi_{i,j}(z)]_{1\leq i,j\leq n}$ is unitary and
$\left[\phi_{i,j}(z)\right]_{1\leq i,j\leq n}\begin{bmatrix}f_{1}(z),f_{2}(z),\cdots,f_{n}(z)\end{bmatrix}^{\rm tr}=\begin{bmatrix}g_{1}(z),g_{2}(z),\cdots,g_{n}(z)\end{bmatrix}^{\rm tr}.$
However, it is not clear whether the functions $\phi_{i,j}$ can be chosen to be Borel measurable. The next theorem provides a positive answer.
Theorem 4.1.
Let $\{f_{i}\}_{1\leq i\leq n}$ and $\{g_{i}\}_{1\leq i\leq n}$ be two families of functions in $L^{2}(\mu)$. Then, the following equation
[…]
has a solution $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$ if and only if $\sum_{k=1}^{n}|f_{k}|^{2}=\sum_{k=1}^{n}|g_{k}|^{2}$ $\mu$-a.e.
Lemma 4.2.
Let $U\in\mathcal{B}\left(L^{2}(\mu)^{(n)}\right)$. Then, $U$ is unitary and commutes with $M_{\mu}^{(n)}$ if and only if $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$ for some $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$.
Proof.
Suppose $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$ for some $[\phi_{i,j}]_{1\leq i,j\leq n}\subset L^{\infty}(\mu)$. Then,
$UU^{*}$ $=\left[\sum_{k=1}^{n}M_{\phi_{i,k}}M_{\phi_{j,k}}^{*}\right]_{1\leq i,j\leq n}=\left[\sum_{k=1}^{n}M_{\phi_{i,k}}M_{\overline{\phi_{j,k}}}\right]_{1\leq i,j\leq n}$
$=\left[\sum_{k=1}^{n}M_{\phi_{i,k}\overline{\phi_{j,k}}}\right]_{1\leq i,j\leq n}=\left[M_{\sum_{k=1}^{n}\phi_{i,k}\overline{\phi_{j,k}}}\right]_{1\leq i,j\leq n}.$
Similarly, $U^{*}U=\left[M_{\sum_{k=1}^{n}\overline{\phi_{k,i}}\phi_{k,j}}\right]_{1\leq i,j\leq n}$. Hence, the operator $U$ is unitary if and only if, for all $1\leq i,j\leq n$,
[…]
where $\delta_{ij}$ denotes the Kronecker delta. That is, $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$. The desired equivalence follows now directly from [8, Page 276, Proposition 6.1 (b)] and [8, Page 279, Corollary 6.9]. ∎
Proof of Theorem 4.1.
In the light of the previous lemma, the equation (4.1) has a solution in $\mathcal{U}(\mu,n)$ if and only if there exists a unitary operator $U$ commuting with $M_{\mu}^{(n)}$ such that $UF=G$, where
$F=\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}\quad\mbox{and}\quad G=\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}.$
By Theorem 2.4, this is true if and only if $\left\|E_{M_{\mu}^{(n)}}(\Delta)F\right\|_{L^{2}(\mu)^{(n)}}=\left\|E_{M_{\mu}^{(n)}}(\Delta)G\right\|_{L^{2}(\mu)^{(n)}}$ for every Borel subset $\Delta\subseteq\mathbb{C}$. Note that
$E_{M_{\mu}^{(n)}}(\Delta)=\bigoplus_{k=1}^{n}E_{M_{\mu}}(\Delta)=\bigoplus_{k=1}^{n}M_{\chi_{\Delta}}$
(see, for instance, [9, Page 49, Proposition 10.4 (a)]). Consequently, (4.1) has a solution in $\mathcal{U}(\mu,n)$ if and only if, for every Borel subset $\Delta\subseteq\mathbb{C}$, we have
$\int\sum_{k=1}^{n}|f_{k}|^{2}\chi_{\Delta}\,{\rm d}\mu$ $=\sum_{k=1}^{n}\int|f_{k}|^{2}\chi_{\Delta}\,{\rm d}\mu=\sum_{k=1}^{n}\left\|E_{M_{\mu}}(\Delta)f_{k}\right\|_{L^{2}(\mu)}^{2}=\left\|E_{M_{\mu}^{(n)}}(\Delta)F\right\|_{L^{2}(\mu)^{(n)}}^{2}$
$=\left\|E_{M_{\mu}^{(n)}}(\Delta)G\right\|_{L^{2}(\mu)^{(n)}}^{2}=\int\sum_{k=1}^{n}|g_{k}|^{2}\chi_{\Delta}\,{\rm d}\mu.$
That is, $\sum_{k=1}^{n}|f_{k}|^{2}=\sum_{k=1}^{n}|g_{k}|^{2}$ $\mu$-a.e. ∎
The measure $\mu$ is called a symmetric measure if $\mu(\Delta)=\mu(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$.
Theorem 4.3.
Assume that $\mu$ is symmetric, and let $\{f_{i}\}_{1\leq i\leq n}$ and $\{g_{i}\}_{1\leq i\leq n}$ be two families of functions in $L^{2}(\mu)$. Then, the following equation
[…]
has a solution $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{SU}(\mu,n)$ if and only if $\sum_{k=1}^{n}|f_{k}(z)|^{2}=\sum_{k=1}^{n}|g_{k}(z)|^{2}$ and $\sum_{k=1}^{n}f_{k}(z)g_{k}(-z)=\sum_{k=1}^{n}f_{k}(-z)g_{k}(z)$ for $\mu$-almost every $z\in\mathbb{C}$.
Lemma 4.4.
Let $U\in{\mathcal{B}}(H)$ be a unitary operator and $J$ a conjugation on $H$. Then, $U=JC$ for some conjugation $C$ on $H$ if and only if $JUJ=U^{*}$.
Proof.
In the light of the facts that $J^{2}=I$ and $\langle JUh,JUk\rangle=\langle Uk,Uh\rangle=\langle k,h\rangle$ for all $h,k\in H$, it is easy to see that
[…]
$\iff(JU)=(JU)^{-1}=U^{*}J^{-1}$
$\iff JUJ=U^{*}.$
Proof of Theorem 4.3.
First, note that the map $J$ defined by $Jf(z)=\overline{f(-z)}$ is a conjugation on $L^{2}(\mu)$. In fact, $J$ is a well-defined conjugate-linear map with $J^{2}=I_{L^{2}(\mu)}$. Additionally, since $\mu$ is symmetric, we have
$\langle Jf,Jg\rangle_{L^{2}(\mu)}=\int\overline{f(-z)}g(-z)\,{\rm d}\mu=\int\overline{f(z)}g(z)\,{\rm d}\mu=\langle g,f\rangle_{L^{2}(\mu)}$
for all $f,g\in L^{2}(\mu)$.
Also, for every $\phi\in L^{\infty}(\mu)$, we have
$JM_{\phi}Jf(z)=JM_{\phi}\overline{f(-z)}=J\phi(z)\overline{f(-z)}=\overline{\phi(-z)}f(z)$
for every $f\in L^{2}(\mu)$. Hence,
(4.3) $JM_{\phi}J=M_{\widetilde{\phi}}.$
In particular, $J^{(n)}M_{\mu}^{(n)}J^{(n)}=-M_{\mu}^{(n)*}$, and $M_{\mu}^{(n)}$ is skew-symmetric since $J^{(n)}$ is a conjugation on $L^{2}(\mu)^{(n)}$.
Claim: An operator $U\in\mathcal{B}\left(L^{2}(\mu)^{(n)}\right)$ is unitary, commutes with $M_{\mu}^{(n)}$ and satisfies $J^{(n)}UJ^{(n)}=U^{*}$ if and only if $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$ for some $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{SU}(\mu,n)$. By Lemma 4.2, it suffices to show that if $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$, then
$J^{(n)}UJ^{(n)}=U^{*}\quad\Longleftrightarrow\quad\phi_{i,j}(z)=\phi_{j,i}(-z)\;\mu\mbox{-a.e. for all }1\leq i,j\leq n.$
In fact, simple matrix calculations show that
$J^{(n)}UJ^{(n)}=U^{*}$ $\quad\Longleftrightarrow\quad\left(JM_{\phi_{i,j}}J\right)_{1\leq i,j\leq n}=\left(M_{\phi_{j,i}}^{*}\right)_{1\leq i,j\leq n}$
$\quad\Longleftrightarrow\quad\left(JM_{\phi_{i,j}}J\right)_{1\leq i,j\leq n}=\left(M_{\overline{\phi_{j,i}}}\right)_{1\leq i,j\leq n}$
[…]
$\quad\Longleftrightarrow\quad\phi_{i,j}(z)=\phi_{j,i}(-z)\;\;\mu\mbox{-a.e. for all }1\leq i,j\leq n.$
This proves the claim.
Now, one can see that the equation (4.2) has a solution in $\mathcal{SU}(\mu,n)$ if and only if, there exists a unitary operator $U$ on $L^{2}(\mu)^{(n)}$ commuting with $M_{\mu}^{(n)}$ such that $J^{(n)}UJ^{(n)}=U^{*}$ and
$U\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}=\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}.$
By Lemma 4.4, this true if and only if there exists a conjugation $C$ on $L^{2}(\mu)^{(n)}$ such that $J^{(n)}C$ commutes with $M_{\mu}^{(n)}$ and
(4.4) $J^{(n)}C\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}=\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}.$
Note that
$J^{(n)}CM_{\mu}^{(n)}=M_{\mu}^{(n)}J^{(n)}C$ $\quad\Longleftrightarrow\quad CM_{\mu}^{(n)}C=J^{(n)}M_{\mu}^{(n)}J^{(n)}$
[…]
Hence, taking into account (4.4), the equation (4.2) has a solution in $\mathcal{SU}(\mu,n)$ if and only if there exists a conjugation $C$ on $L^{2}(\mu)^{(n)}$ such that $CM_{\mu}^{(n)}C=-M_{\mu}^{(n)*}$ and
$C\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}=J^{(n)}\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}=\begin{bmatrix}\widetilde{g_{1}},\widetilde{g_{2}},\cdots,\widetilde{g_{n}}\end{bmatrix}^{\rm tr}.$
For convenience, put $F=\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}$ and $\widetilde{G}=\begin{bmatrix}\widetilde{g_{1}},\widetilde{g_{2}},\cdots,\widetilde{g_{n}}\end{bmatrix}^{\rm tr}$. Then, by Theorem 3.2, the equation (4.2) has a solution in $\mathcal{SU}(\mu,n)$ if and only if
$\left\langle E_{M_{\mu}^{(n)}}(\Delta)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle E_{M_{\mu}^{(n)}}(-\Delta)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
and
$\left\langle E_{M_{\mu}^{(n)}}(\Delta)F,F\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle E_{M_{\mu}^{(n)}}(-\Delta)\widetilde{G},\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
for every Borel subset $\Delta\subseteq\mathbb{C}$. That is,
$\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{\Delta}}\right)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{-\Delta}}\right)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
and
$\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{\Delta}}\right)F,F\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{-\Delta}}\right)\widetilde{G},\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
for every Borel subset $\Delta\subseteq\mathbb{C}$; which means that
$\int\sum_{i=1}^{n}f_{i}(z)g_{i}(-z)\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}f_{i}(z)g_{i}(-z)\chi_{-\Delta}\,{\rm d}\mu$
and
$\int\sum_{i=1}^{n}|f_{i}(z)|^{2}\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}|g_{i}(-z)|^{2}\chi_{-\Delta}\,{\rm d}\mu$
for every Borel subset $\Delta\subseteq\mathbb{C}$. Therefore, by symmetry of $\mu$, for every Borel subset $\Delta\subseteq\mathbb{C}$,
$\int\sum_{i=1}^{n}f_{i}(z)g_{i}(-z)\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}f_{i}(-z)g_{i}(z)\chi_{\Delta}\,{\rm d}\mu$
and
$\int\sum_{i=1}^{n}|f_{i}(z)|^{2}\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}|g_{i}(z)|^{2}\chi_{\Delta}\,{\rm d}\mu;$
which lead to the desired conditions. ∎

### v3

@@ 3.1. Main results
$Vp(N_{k_{1}},N_{k_{1}}^{*},\dots,N_{k_{n}},N_{k_{n}}^{*})y_{i}$ $=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}Vy_{i}$
$=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}x_{i}.$
Put $\mathcal{L}_{0}=L_{\mathcal{N},X\cup Y}$. Then, the above equalities imply that $V\mathcal{L}_{0}=\mathcal{L}_{0}$. Since $V$ is anti-unitary, we get $V\mathcal{L}_{1}=V\mathcal{L}_{0}^{\perp}=\mathcal{L}_{0}^{\perp}=\mathcal{L}_{1}$.
Let $V_{1}$ denote the restriction of $V$ to $\mathcal{L}_{1}$. Then, $V_{1}$ is anti-unitary on $\mathcal{L}_{1}$, and the relations $VN_{k}V^{-1}=-N_{k}^{*}$, for every $k\in K$, imply $V_{1}N_{k|\mathcal{L}_{1}}V_{1}^{-1}=-\left(N_{k|\mathcal{L}_{1}}\right)^{*}$. This proves the first part.
Now, by Lemma 2.15, for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $k\in K$, we have $VE_{N_{k}}(-\Delta)=E_{N_{k}}(\Delta)V$. The remainder of the proof is now similar to (2.8) and (2.9).
(iii)$\implies$(i). For each $k\in K$, put $M_{k}=-N_{k}$. Let $K_{0},K_{1}\subseteq K$ be finite subsets, and let $\{\Delta_{k}\}_{k\in K_{0}}$ and $\{D_{k}\}_{k\in K_{1}}$ be Borel subsets of $\mathbb{C}$. Then, for all $i,j\in I$, we have
$\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)x_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{-N_{l}}(D_{l})\right)x_{j}\right\rangle$
$=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{N_{l}}(-D_{l})\right)x_{j}\right\rangle$
$=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(-\Delta_{k})\right)x_{j},\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{i}\right\rangle$
$=\left\langle\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)x_{j},\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{i}\right\rangle.$
Similarly,
$\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)y_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)x_{j},\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)y_{i}\right\rangle.$
It follows by Theorem 2.1 that there exists a conjugation $C_{0}$ on $\mathcal{L}_{0}:=L_{\mathcal{N},X\cup Y}$ such that $C_{0}{N_{k}}_{|\mathcal{L}_{0}}C_{0}=-\left({N_{k}}_{|\mathcal{L}_{0}}\right)^{*}$ and $C_{0}x_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
Now, since by hypothesis $V_{1}{N_{k}}_{|\mathcal{L}_{1}}V_{1}^{-1}=-\left({N_{k}}_{|\mathcal{L}_{1}}\right)^{*}$ for every $k\in K$, Lemma 3.7 implies that there exists a conjugation $C_{1}$ on $\mathcal{L}_{1}$ so that $C_{1}{N_{k}}_{|\mathcal{L}_{1}}C_{1}=-\left({N_{k}}_{|\mathcal{L}_{1}}\right)^{*}$ for every $k\in K$. Therefore, $C=C_{0}\oplus C_{1}$ is a conjugation on $\mathcal{H}=\mathcal{L}_{0}\oplus\mathcal{L}_{1}$ that satisfies all required equalities. ∎
@@
(ii)$\implies$(i). Let $\mathcal{H}_{0}=L_{\{N\},X\cup Y}$ and $\mathcal{H}_{1}=\mathcal{H}_{0}^{\perp}$, and denote by $N_{0}$ and $N_{1}$ the restrictions of $N$ to $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, respectively. In the light of Theorem 3.2 and Lemma 3.7, we only need to show that $N_{1}$ is skew-symmetric. This follows immediately from Theorem 3.10 and the facts that $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$ and $N=N_{0}\oplus N_{1}$ is skew-symmetric.
@@
$\mathcal{H}_{0}=\bigvee\left(\sum_{i\in I}L_{\{N\},\{x_{i}\}}+L_{\{N\},\{y_{i}\}}\right)=\bigvee\left(\sum_{i\in I}L_{\{N_{0}\},\{x_{i}\}}+L_{\{N_{0}\},\{y_{i}\}}\right)$
@@
and the restrictions of $N_{0}$ to $L_{\{N_{0}\},\{x_{i}\}}$ and $L_{\{N_{0}\},\{y_{i}\}}$ are star-cyclic, Lemma 3.11 implies that $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$. This concludes the proof. ∎